OpenStudy (anonymous):

Determine which equation below has a solution of x = 2.A. –x + 9 = 14 – 3x B. –4x + 3 = 8x – 21 C. 3x – 6 = 2x + 4 D. 18x – 32 = –23x – 6

OpenStudy (anonymous):

for each option, all you need to do is solve for x. so for A, -x+9 = 14, you just subtract 9 from both sides and you get -x = -5. multiply both sides by -1 to get x = 5. A is not your answer. do a similar thing for each other option until A is = to 2

OpenStudy (anonymous):

oh sorry. read A wrong XD it's -x + 9 = 14 - 3x sorry my bad. so instead add the 3x to both sides because it's the larger value of x and you get 2x, again, subtract 9 from both sides and you get 5. divide both sides by 2 and you get 5/2 so still A is not your answer

OpenStudy (anonymous):

hope that helps